### v1

@@ 3.5 Summary: From Observations to Design Principles
These failure modes reinforce one another. Controller brittleness consumes context that could otherwise support broader exploration; incomplete exploration leads to missed evidence; and rigid reasoning states prevent recovery when initial assumptions prove incorrect. The following section formalizes an architecture that addresses these deficiencies through the Explanations-over-Graphs framework.
@@ 4 Methodology
Figure 5: Semantic Belief Propagation architecture, illustrated with the worked example (§4.5). Left: The topology graph $G$ is incomplete—the $S_{2}\to S_{3}$ dependency is not registered; additional services ($S_{5}$$S_{7}$) are uninvolved (healthy). An alert on $S_{2}$ seeds the ActiveSet. Center: The active inference loop: pop from ActiveSet, fetch context, classify via policy, update state. New discoveries or belief changes add nodes back to ActiveSet. Right: The output explanatory graph $G_{S}$ with the discovered edge (dashed) and final beliefs. Origin nodes use double circles; symptom nodes use hatched fill. The frontier $\mathcal{F}=\{S_{1}\}$ identifies the flash sale as the true origin.
We propose Semantic Belief Propagation (SBP), an architecture designed to address the structural failures of ReAct agentic loops identified in §3.5: unguided exploration (O3, I1), controller brittleness (O4, I2), and path-dependent reasoning (O5, I3). We implement this architecture as the EoG (Explanations over Graph) agent, which performs distributed causal inference over operational graphs, in contrast with existing monolithic or hierarchical approaches typically relying on the LLM to simultaneously manage exploration, reasoning, and global state. This coupling makes the resulting existing systems fragile: a single planning error or context overflow causes the entire diagnosis to fail.
SBP decomposes this loop into a distributed message-passing algorithm inspired by classical belief propagation Pearl (1988), enforcing a strict separation of concerns:
  • State (The Explanatory Graph): Instead of relying on an ephemeral context window, we maintain a persistent Explanatory Graph $G_{S}$. This structure acts as explicit external memory, storing the evolving belief state (e.g., $S_{1}$ is an Origin, $S_{2}$ is a Symptom) separate from the static system topology—addressing I3.
  • Decision (The Policy): We restrict the LLM to a stateless Abductive Policy $\pi_{\text{abd}}$. By removing the burden of global planning, the policy can focus entirely on local inference—mapping bounded evidence (logs, neighbor messages) to causal roles—addressing I2.
  • Control (The Algorithm): We replace the unstable “LLM-as-Planner” with a Deterministic Controller. This symbolic engine offloads housekeeping duties (queue management, message routing, termination), enforcing a rigorous exploration strategy that prevents the agent from “getting lost” or abandoning the plan, addressing I1.
This architecture enables non-monotonic reasoning: as new evidence propagates (e.g., a downstream service reports it is a symptom), upstream nodes can revise their initial beliefs, correcting early misdiagnoses without restarting the investigation.
Handling Inconclusive Results.
The agent may not always produce a definitive diagnosis. Two failure modes are possible: (1) evidence absence—the relevant observability data does not exist or is not retrievable via the CxC (e.g., logs were not retained, metrics not instrumented); and (2) evidence blindness—the policy $\pi_{\text{abd}}$ lacks the domain knowledge to recognize available evidence as diagnostically relevant. In either case, nodes are labeled Defer. If no Origin is identified at termination (i.e., $\mathcal{F}=\emptyset$), the Controller falls back to a best-effort ranking of candidate nodes based on accumulated evidence, explicitly signaling uncertainty to the operator.
@@ 4.1 Architecture Overview
The architecture comprises three components (Figure 5):
The Deterministic Controller.
A symbolic engine that maintains the global Ledger, manages the ActiveSet (nodes to investigate), and routes messages. It executes the algorithm logic (termination, damping) without performing semantic reasoning. This separation ensures safety and reproducibility: the agent cannot execute arbitrary code or tools; it can only request data via the Context Contract, preventing harmful side effects. It also provides operational efficiency: graph traversal, message routing, and bookkeeping require no LLM calls, bounding inference costs and API utilization to $O(|V_{S}|)$ policy invocations.
The Context Contract (CxC).
A strict interface that bounds and focuses the information provided to the policy, drawing on principles of active context management Li et al. (2025); Zhang et al. (2025a). When investigating node $v$, the CxC fetches a Context Packet containing:
  1. 1.
    Local observability data (logs, metrics) within the incident time window.
  2. 2.
    Topological context (immediate neighbors).
  3. 3.
    Incoming messages from neighbors (the Inbox).
@@
The Abductive Policy ($\pi_{\text{abd}}$).
A stateless, LLM-parameterized policy that performs local abductive inference. Given a bounded Context Packet for node $v$, $\pi_{\text{abd}}$ produces: (i) a local belief $B_{v}$, (ii) propagation claims $\mathcal{P}_{v}$, and (iii) a set of candidate entities $\mathcal{N}_{next}$ for further investigation. While the Controller governs execution and termination, $\pi_{\text{abd}}$ determines the epistemic trajectory of the search. We use “policy” in the decision-theoretic sense: a mapping from local observations to epistemic decisions (belief assignment and hypothesis selection), not as an online reward-optimizing controller.
@@ 4.2 State Formalism
We distinguish between three levels of state:
1. The Global Ledger ($\mathcal{L}$).
The immutable history of all reasoning steps, serving as an external scratchpad Nye et al. (2021) that enables replay and debugging.
2. Local Belief ($B_{v}$).
The current assessment of a node $v$, derived from the latest policy execution: $B_{v}=(L_{v},\mathcal{E}_{v})$ where:
@@
3. Messages ($m_{u\to v}$).
A mechanism to transport beliefs. A message contains a snapshot of $B_{u}$ at time $t$. When node $v$ processes its Inbox, it reads these messages to reason about upstream dependencies (the Local Markov Property).
4.3 Algorithm: Semantic Belief Propagation
We present a unified algorithm that interleaves exploration (finding new nodes) and propagation (refining beliefs). The system is implemented as an actor model, allowing nodes to process messages concurrently.
@@ 4.3.1 Initialization (Bootstrap)
  • Define time window $W$ based on incident timestamp.
  • Initialize $\text{ActiveSet}\leftarrow A$ (or other domain heuristics).
  • Initialize $V_{S}\leftarrow\emptyset,E_{S}\leftarrow\emptyset$.
  • Set $\text{Budget}\leftarrow\text{max number of investigation hopes}$
@@ 4.3.2 Active Inference Loop
6:  $\text{Packet}\leftarrow\text{CxC.Fetch}(v,W,\text{Inbox})$
7:  // 2. Execute Policy
8:  $(B_{v}^{new},\mathcal{P}_{v},\mathcal{N}_{next})\leftarrow\pi_{\text{abd}}(\text{Packet})$
9:  // 3. Damping (Heuristic)
10:  if $\textsc{Flips}(v)>k_{\text{thresh}}$ then
11:   $B_{v}^{new}.L\leftarrow\textsc{Defer}$
12:  end if
13:  // 4. Update State
14:  $\text{Ledger}.\text{append}(v,B_{v}^{new})$
15:  for $(u\to v)\in\mathcal{P}_{v}$ do
16:   $E_{S}.\text{add}(u\to v)$
17:   if $u\notin V_{S}$ then
18:     $\text{ActiveSet}.\text{add}(u)$ $\triangleright$ Exploration
19:   end if
20:  end for
21:  // 5. Propagation
22:  if $B_{v}^{new}\neq B_{v}^{old}$ then
23:   $\text{Broadcast}(B_{v}^{new}\text{ to }\mathcal{N}(v))$
24:   $\text{ActiveSet}.\text{add}(\mathcal{N}(v))$ $\triangleright$ Re-activate neighbors
25:  end if
26:  // 6. Propose (Optional)
27:  for $u\in\mathcal{N}_{next}$ do
28:   if $u\notin V_{S}$ then
29:     $\text{ActiveSet}.\text{add}(u)$
30:   end if
31:  end for
32:  $\text{Budget}$-- $\triangleright$ Decrement Budget
33:end while
34:// Post-processing: Compute Frontier
@@
The loop terminates when no more messages are pending (fixed point) or budget is exhausted. The Frontier $\mathcal{F}$ is computed after convergence as a post-processing step over the final belief state.
4.4 Properties and Correctness
Minimality and Correctness.
The algorithm seeks a Frontier $\mathcal{F}$ of irreducible origins.
$\mathcal{F}=\{v\in V_{S}:$ $L_{v}=\textsc{Origin}\land\nexists u\in V_{S}$ (1)
$\text{s.t. }L_{u}=\textsc{Origin}\land u\rightsquigarrow_{E_{S}}v\}$
By definition, $\mathcal{F}$ is minimal relative to the discovered graph $G_{S}$. If no node is labeled Origin (i.e., $\mathcal{F}=\emptyset$), the finalize phase falls back to best-effort ranking of candidates based on accumulated evidence. The correctness relies on three assumptions:
  1. 1.
    Completeness: Relevant data exists and is retrievable via CxC.
  2. 2.
  3. 3.
    Connectivity: Origins are reachable from alerts via the topology or discovered edges.
If these hold, the message passing ensures that evidence from origins propagates to symptoms, correcting any initial misclassifications caused by exploration order.
Termination.
4.5 Worked IT Example: Cascading Belief Revision
@@
The diagnostic trap.
$S_{4}$’s Out Of Memory (OOM) error is strong local evidence of being the origin—memory exhaustion is a classic root cause. Without belief messages, the algorithm would terminate with $S_{4}$ as origin, leading operators to restart $S_{4}$ (which doesn’t address the upstream traffic pattern). This example shows both edge discovery (finding the hidden $S_{2}\to S_{3}$ dependency) and belief revision (correcting the $S_{4}$ misdiagnosis).
@@
Figure 6: Edge discovery and belief revision. Row 1: The hidden edge $S_{3}\to S_{2}$ (dashed) is discovered via HTTP 500 traces—it was not in the topology graph. $S_{4}$’s OOM is the sole origin ($\mathcal{F}=\{S_{4}\}$), but traffic analysis reveals elevated load; $S_{4}$ revises and broadcasts. Row 2: Messages cascade upstream; $S_{1}$ (flash sale) is identified as the true origin. Origin nodes use double circles; symptom nodes use hatched fill.
@@
    $t=3$, Process $S_{4}$: Logs show OutOfMemoryError. Strong local evidence of being the origin—memory exhaustion is a classic root cause. $B_{S_{4}}=\textsc{Origin}$.
    At this point: $\mathcal{F}=\{S_{4}\}$. Single origin, plausible diagnosis. Without belief messages, we would stop here and blame $S_{4}$’s memory issue.
@@
    $t=5$, $S_{3}$ receives message (Alg. Step 5: Propagation continues): Inbox contains $B_{S_{4}}$: “I failed from your traffic.” $S_{3}$ checks its own incoming traffic—+25% from $S_{2}$. Propagates message upstream. Broadcasts to $S_{2}$.
@@
    $t=6$, $S_{2}$ receives message: Inbox contains $B_{S_{3}}$: “The cascade started from your traffic.” $S_{2}$ checks incoming—+20% from $S_{1}$. Proposes $S_{1}$.
@@
Key insight.
[…]
@@ 5 Implementation & Results
5.1 Implementation
We implement the EoG agent and a baseline ReAct agent on top of the Codex agent framework. [^1] Available at https://github.com/openai/codex The implementation comprises approximately 5,000 lines of Rust code. The Deterministic Controller manages the ActiveSet and Ledger with read-write locks, enabling concurrent node processing while ensuring atomic belief updates. The actor-model implementation supports asynchronous execution without corrupting global state even when messages arrive out of order. Observability data is retrieved through the Model Context Protocol (MCP), a standardized interface for LLM tool use. We implement domain-specific MCP servers exposing the operational graph and observability signals—including distributed traces Kaldor et al. (2017)—via the tools in Table 1.
@@ 5.1 Implementation
event_analysis Platform events in time window
@@
spec_change_analysis Configuration/deployment diffs
Table 1: MCP tools for SRE domain context discovery. EoG uses only alert_summary (bootstrap) and get_context_contract (exploration); the ReAct baseline has access to all tools, shell utilities, and can write Python code for data analysis.
5.2 Experiments
We seek to answer the following research questions:
RQ1.
Does the EoG agent improve majority@$k$ and pass@$k$ over baselines?
RQ2.
Can ReAct agents achieve similar results when the EoG algorithm is explained in the prompt?
RQ3.
How much does SBP improve accuracy beyond the Deterministic Controller alone?
Setup.
We evaluate our methodology on ITBench (Jha et al., 2025), a benchmark of realistic SRE diagnostic tasks on Kubernetes, complementing other agentic task automation benchmarks Shen et al. (2024); Patel et al. (2025); Qiao et al. (2025). We use 35 scenarios spanning misconfigurations, resource exhaustion, and cascading failures. Each scenario is repeated 3 times to measure accuracy and consistency. All experiments use high reasoning effort settings (or equivalent) for each model. We compare EoG against a ReAct baseline; notably, ReAct has access to more tools (all MCP tools plus shell utilities) while EoG uses only two MCP calls per entity, yet controls the investigation structure algorithmically.
RQ1: EoG vs ReAct.
Table 2 compares EoG against ReAct baselines. The results directly address the reliability gap identified in §3.1: while ReAct exhibits large variance between pass@3 and majority@3 (indicating stochastic success), EoG narrows this gap substantially. For example, GPT-5.1 under ReAct achieves only 22.9% pass@3 and 8.6% majority@3 F1 (RC Entity); under EoG, these rise to 88.9% and 86.1% respectively—a near-closed reliability gap.
[…]
We note that the ReAct baseline was provided access to operational topology via tools and the application service architecture in the system prompt—the performance gap is not due to missing structural information. Additionally, ablation experiments removing distributed traces as a data source showed no statistically significant change in results; this is consistent with the redundancy typical of observability data, where the same diagnostic evidence can often be inferred from multiple signals (logs, metrics, events).
Table 2: RQ1: EoG vs ReAct on ITBench (35 scenarios, 3 runs).
@@ 5.2 Experiments
  • Best overall shown in bold; other EoG wins in light blue. “–” indicates ReAct baseline could not be run due to non-trivial failure modes with the Codex framework. We omit diagnosis time as it varies due to provider rate limits, reasoning token overhead, and concurrency. Token counts are averaged per trial per scenario and exclude reasoning tokens as providers do not consistently report them.
RQ2: Can ReAct be prompted to follow EoG?
We investigate whether the reliability gains of EoG can be achieved simply by instructing a ReAct agent to follow the EoG algorithm (i.e., “explore neighbors, update beliefs, propagate messages”) in its system prompt, without the external Deterministic Controller. We select a diverse subset of high-performing models due to budget constraints. Table 3 shows the results.
While ReAct+Prompting achieves non-trivial pass@3 recall (40–71%), its majority@3 recall collapses to near zero (0–2.9%). This indicates extreme instability: the agent may occasionally stumble upon the correct diagnosis but cannot reproduce it reliably. The cognitive load of simultaneously managing exploration state, message passing, and causal reasoning overwhelms the model’s instruction-following capacity. This confirms that algorithmic enforcement via a symbolic controller is essential; the EoG protocol cannot be implemented reliably as a prompt-only policy.
Table 3: RQ2: ReAct prompted with EoG algorithm (Recall).
@@
RQ3: SBP Ablation.
Table 4 isolates the contribution of the two architectural components: the Deterministic Controller (graph-guided traversal) and SBP (non-monotonic revision). EoG-SBP uses only the Controller—entities are classified once without revision. Full EoG adds SBP, allowing upstream beliefs to be revised when downstream evidence contradicts initial assessments.
The results show that structured traversal alone (EoG-SBP) captures significant improvement over ReAct, confirming that the exploration pathology (§3.2) is a dominant failure mode. SBP provides substantial additional gains of 7–64% across models, with the largest improvements on weaker models (GPT-OSS-20B: +64.3%) where early misclassifications are more common. This validates the order sensitivity findings from §3.4: while oracle reordering showed 14–70% potential gains from optimal sequencing, SBP recovers a significant portion of this by enabling post-hoc correction without requiring oracle knowledge.
Table 4: RQ3: SBP ablation (maj@3 Recall, RC Entity).

### v2

@@ 3.3 Controller Failures
These failure modes reinforce one another. Controller brittleness consumes context that could otherwise support broader exploration. Incomplete exploration leads to missed evidence. Path-dependent reasoning prevents recovery when initial assumptions prove incorrect.
@@ 4 Methodology
Figure 3: EoG agent architecture. (1) Topology $G$ with missing edge; alert seeds ActiveSet. (2) Loop: pop node, fetch context (CxC), run policy $\pi_{\text{abd}}$, update $G_{S}$. Expand adds new neighbors; Propagate reactivates on belief change. (3) Output $G_{S}$ with discovered edge and frontier $\mathcal{F}=\{S_{1}\}$.
We present EoG (Explanations over Graphs), an agent that traverses the dependency graph $G$ to identify the frontier $\mathcal{F}$: the minimal set of origin nodes whose faults explain all observed anomalies (§2). EoG builds an Explanatory Graph $G_{S}$ where each visited node carries a belief: its inferred explanatory role (Origin, Symptom, or Healthy) and supporting evidence. The agent addresses the limitations of ReAct (§3) through two design principles:
  1. 1.
    Separation of decision and control. A Deterministic Controller handles graph traversal, state management, and tool orchestration, while a stateless Abductive Policy $\pi_{\text{abd}}$ performs local inference over raw evidence (e.g., logs, metrics retrieved via tools) and neighbor messages for each node $v\in V_{S}$. This separation eliminates controller failures (§3.3), ensures auditability, and enforces systematic graph exploration.
  2. 2.
    Belief revision via Semantic Belief Propagation (SBP). Initial beliefs may be incorrect due to incomplete information. For instance, a node with strong failure symptoms (e.g., out-of-memory) may be a victim of upstream problems (e.g., increased load). When new evidence changes a node’s belief, it broadcasts to its neighbors, triggering re-evaluation. This message-passing mechanism, inspired by belief propagation Pearl (1988), addresses path dependence (§2), by correcting early misdiagnoses as evidence propagates, without restarting the investigation.
@@ 4.1 Architecture Overview
The architecture comprises three modules (Figure 3):
The Deterministic Controller. A symbolic engine that maintains all persistent state: the Explanatory Graph $G_{S}=(V_{S},E_{S})$ with node beliefs, the Global Ledger $\mathcal{L}$ (immutable history of reasoning steps), the ActiveSet (nodes pending investigation), and message queues, which execute graph traversal, message routing, and termination logic without semantic reasoning. This separation ensures safety and reproducibility: the agent cannot execute arbitrary code; it can only request data via the Context Contract (a specialized tool described below). It also provides operational efficiency: all bookkeeping requires no LLM calls, bounding inference costs to policy invocations.
The Context Contract (CxC). A strict interface that bounds and focuses the information provided to the policy, drawing on principles of active context management Li et al. (2025); Zhang et al. (2025a). The CxC is implemented as an MCP tool anthropic2024mcp with a simple interface get_context(entity_id, start_time, end_time), abstracting away domain-specific details. When investigating node $v$, the CxC returns local evidence on that entity (e.g., logs, metrics, traces in IT systems) and topological context (immediate neighbors). The Controller bundles this with the Inbox to form the Context Packet.
@@
The Abductive Policy ($\pi_{\text{abd}}$). A stateless, LLM-parameterized policy that performs local abductive inference. Given a bounded Context Packet for a node $v$, $\pi_{\text{abd}}$ produces: (i) a local belief $B_{v}$, (ii) propagation claims $\mathcal{P}_{v}$, and (iii) a set of candidate entities $\mathcal{N}_{next}$ for further investigation. While the Controller governs execution and termination, $\pi_{\text{abd}}$ determines the epistemic trajectory of the search. We use “policy” in the decision-theoretic sense: a mapping from local observations to belief assignment.
@@ 4.2 State Formalism
EoG maintains following state variables:
1. The Global Ledger ($\mathcal{L}$): an append-only log recording each $(v,B_{v},t)$ tuple tracking which node was evaluated, its resulting belief, and when. This scratchpad Nye et al. (2021) enables replay, debugging, and auditability.
2. Local Belief ($B_{v}$): stored per node in $G_{S}$, updated after each policy execution. $B_{v}=(L_{v},\mathcal{E}_{v})$ where:
@@
3. Messages ($m_{u\to v}$): stored in per-node Inboxes. A message contains the sender’s belief $B_{u}$ at the time of broadcast. When node $v$ is evaluated, the policy reads its Inbox to reason about neighbor states, exploiting the Local Markov Property Pearl (1988).
4.3 Algorithm
We present the algorithmic steps of the EoG agent, which interleaves exploration (discovery of new nodes) and propagation ( belief refinement).
@@ 4.3.1 Initialization (Bootstrap)
  • Define time window $W$ (e.g., incident start/end time).
  • Initialize:
    • $\text{ActiveSet}\leftarrow V_{\mathcal{O}}$ (e.g., alerting entities)
    • $V_{S}\leftarrow\emptyset,E_{S}\leftarrow\emptyset$
    • $\text{Budget}\leftarrow\text{max investigation hops}$
@@ 4.3.2 Active Inference Loop
6:  $\text{LocalContext}\leftarrow\text{CxC.Fetch}(v,W)$
7:  $\text{Packet}\leftarrow(\text{LocalContext},\text{Inbox})$
8:  // 2. Execute Policy
9:  $(B_{v}^{new},\mathcal{P}_{v},\mathcal{N}_{next})\leftarrow\pi_{\text{abd}}(\text{Packet})$
10:  // 3. Damping (Heuristic)
11:  if $\textsc{Flips}(v)>k_{\text{thresh}}$ then
12:   $B_{v}^{new}.L\leftarrow\textsc{Defer}$
13:  end if
14:  // 4. Update State
15:  $\text{Ledger}.\text{append}(v,B_{v}^{new})$
16:  for $(u\to v)\in\mathcal{P}_{v}$ do
17:   $E_{S}.\text{add}(u\to v)$
18:   if $u\notin V_{S}$ then
19:     $\text{ActiveSet}.\text{add}(u)$ $\triangleright$ Exploration
20:   end if
21:  end for
22:  // 5. Propagation
23:  if $B_{v}^{new}\neq B_{v}^{old}$ then
24:   $\text{Broadcast}(B_{v}^{new}\text{ to }\mathcal{N}(v))$
25:   $\text{ActiveSet}.\text{add}(\mathcal{N}(v))$ $\triangleright$ Re-activate neighbors
26:  end if
27:  // 6. Propose (Optional)
28:  for $u\in\mathcal{N}_{next}$ do
29:   if $u\notin V_{S}$ then
30:     $\text{ActiveSet}.\text{add}(u)$
31:   end if
32:  end for
33:  $\text{Budget}$-- $\triangleright$ Decrement Budget
34:end while
35:// Post-processing: Compute Frontier
@@
37:// Finalize: Generate Report
@@
The loop terminates when no more messages are pending (fixed point) or the budget is exhausted. The Frontier $\mathcal{F}$ is computed after convergence, and a finalize policy $\pi_{\text{finalize}}$ synthesizes findings into a structured report (Appendix B.7). Additional details on correctness properties, termination guarantees, and handling of inconclusive results are provided in Appendix B.9.
4.4 Example scenario.
@@
The diagnostic trap. $S_{4}$’s Out Of Memory (OOM) error is strong local evidence of being the origin. Without belief messages, the algorithm would terminate with $S_{4}$ as origin, leading operators to restart $S_{4}$ (without addressing the upstream traffic pattern). This example shows both edge discovery (finding the hidden $S_{2}\to S_{3}$ dependency) and belief revision (correcting the $S_{4}$ misdiagnosis).
[…]
Figure 4: Edge discovery and belief revision sequence. (1) $S_{2}\to S_{3}$ discovered. (2) $S_{4}$ revises belief due to high load. (3) $S_{3}$ propagates message. (4) $S_{1}$ identified as root cause.
@@
    $t=3$, Process $S_{4}$: Logs show OutOfMemoryError. Strong local evidence of being the origin points to memory exhaustion, that is a classic root cause.
    At this point: $\mathcal{F}=\{S_{4}\}$. Single origin, plausible diagnosis. Without belief messages, this would be the final step of the process pointing to $S_{4}$’s memory issue.
@@
    $t=5$, $S_{3}$ receives the message (Alg. Step 5: Propagation continues): Inbox contains $B_{S_{4}}$: “I failed from your traffic.” $S_{3}$ checks its own incoming traffic: +25% from $S_{2}$. Propagates message upstream. Broadcasts to $S_{2}$.
@@
    $t=6$, $S_{2}$ receives the message: Inbox contains $B_{S_{3}}$: “The cascade started from your traffic.” $S_{2}$ checks incoming: +20% from $S_{1}$. Proposes $S_{1}$.
@@
Key insight. This example demonstrates: (1) structure learning: the hidden from topology graph $S_{2}\to S_{3}$ edge was discovered from observability data Havrilla et al. (2025); and (2) belief revision: $S_{4}$’s plausible OOM diagnosis was corrected via cascading messages. Without SBP, operators would restart $S_{4}$, providing temporary relief until memory fills again. The upstream traffic pattern is never addressed.
@@ 5 Implementation & Results
Table 1: MCP tools for SRE domain.
@@
event_analysis Events in time window
@@
spec_change_analysis Config/deployment diffs
  • The ReAct agent has access to all tools, shell utilities, and can write Python code for data analysis. The EoG agent uses alert_summary, spec_change_analysis, and event_analysis to populate the initial ActiveSet, then invokes get_context_contract during entity visits where abductive policy is executed.
Table 2: EoG vs ReAct on ITBench (35 scenarios, 3 runs).
@@
  • Best overall shown in bold; other EoG wins in light blue. Token counts are averaged per trial per scenario and exclude reasoning tokens as providers do not consistently report them. ReACT with GPT-OSS-20B, Mistral Large and Minimax M2.1 could not be run due to Codex failures.
Table 3: ReAct prompted with EoG algorithm (Recall).
@@
Table 4: SBP ablation (maj@3 Recall, RC Entity).
@@
Implementation.
We implement the EoG agent and a baseline ReAct agent on the Codex agent framework, [^3] Available at https://github.com/openai/codex and evaluate on abductive reasoning tasks in ITBench (§3). The implementation comprises ~5,000 lines of Rust code. The ReAct baseline incorporates best practices from prior work on LLM-based diagnostics Roy et al. (2024), achieving higher performance than published baselines on ITBench; implementation details are provided in Appendices A and B.
We implement domain-specific MCP servers exposing the operational graph and evidence data (e.g., logs, traces, and metrics in ITBench), including distributed traces Kaldor et al. (2017), via the tools in Table 1.
Experiments.
35 ITBench scenarios spanning misconfigurations, resource exhaustion, and cascading failures, are used for evaluation, with 3 runs per scenario to measure both accuracy and consistency. All experiments use high reasoning effort settings. Notably, ReAct has access to more tools (all MCP tools plus shell utilities and Python code execution) while EoG uses only the Context Contract per entity, and controls the investigation structure algorithmically.
We compute two metrics: RC Entity: a correctness of identified root cause entity (e.g., otel-demo/Pod/cartservice-xyz), and RC Reasoning: a correctness of explanation of the causal mechanism (e.g., “Pod crashed due to incorrect OS architecture in container image”). More details in  D.
EoG vs ReAct. Table 2 compares EoG against ReAct baselines. The results directly address the reliability gap identified in §3.1: while ReAct exhibits a large variance between pass@3 and majority@3 (indicating stochastic success), EoG substantially narrows this gap. For example, GPT-5.1 under ReAct achieves only 22.9% pass@3 and 8.6% majority@3 F1 (RC Entity); under EoG, these rise to 88.9% and 86.1% respectively, near-closing the reliability gap.
Additionally, ablation experiments of removing distributed traces as a data source showed no statistically significant change in results; this is consistent with the redundancy typical of observability data, where the same diagnostic evidence can often be inferred from multiple signals (logs, metrics, events).
Prompting Ablation. We quantify the efficacy of a ReAct agent prompted to follow the EoG algorithm (§4.3) without the external Deterministic Controller. Table 3 shows that majority@3 collapses to near zero despite non-trivial pass@3, indicating that the cognitive load of simultaneously managing exploration state, message passing, and causal reasoning overwhelms instruction-following capacity.
SBP Ablation. Table 4 isolates the contribution of SBP: structured traversal alone captures most gains over ReAct, while SBP provides 7–64% additional improvement by enabling post-hoc belief correction (§2), with the largest gains on weaker models where early misclassifications are more common.